Children’s ABO blood types are dependent on their biological parents’. ABO blood type; each parent provides one allele from which a fetus from its ABO blood type. Here is a list of all possible alleles received and their corresponding ABO blood type.

Alleles

AA

AO

AB

BB

BO

OO

Blood type

A

A

AB

B

B

O

If a baby was born to a couple with AO for one parent and AB for the other, what is the probability that this child’s ABO blood type is

  1. A.               b) B              c) AB                d) O
Answer #1

a) probability that this child’s ABO blood type is =P(A Allele from first parent and A allele from 2nd parent)+P(O allele from first parent and A allele from 2nd) =(1/2)*(1/2)+(1/2)*(1/2) =1/2

b)

P(B) =P(O allele from 1st and B from 2nd )=(1/2)*(1/2) =1/4

c)

P(AB) =P(A allele from 1st and B allele from 2nd )=(1/2)*(!/2 )=1/4

d)P(O) =P(O allele from 1st and O allele from 2nd) =(1/2)*0 =0